OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations a.has no solution. b.has one solution. c.is coincident.

OpenStudy (vijay):

2 and -1

OpenStudy (saifoo.khan):

b.has one solution.

OpenStudy (vijay):

:D

OpenStudy (anonymous):

If we multiply an equation by a constant, we do not change the equation. Multiply E1 by 2 and E2 by -3. Then you have: 6x + 4y = 8 -6x + 3y = -15 Add together to eliminate x. 7y = -7 y = -1 Plug in for y in either equation. 6x + 4(-1) = 8 6x - 4 = 8 6x = 12 x = 2 Since there are definite answers for both variables, there is only one solution. If this helps, please leave a medal. Check out my website, where there are free math videos at http://www.tutorsean.net